# EXAdam: The Power of Adaptive Cross-Moments

### v1

@@ Abstract
This paper introduces EXAdam (EXtended Adam), a novel optimization algorithm that builds upon the widely-used Adam [1] optimizer. EXAdam incorporates three key enhancements: (1) new debiasing terms for improved moment estimation, (2) a gradient-based acceleration mechanism for increased responsiveness to the current loss landscape, and (3) a dynamic step size formula that allows for continuous growth of the learning rate throughout training. These innovations work synergistically to address limitations of the original Adam algorithm, potentially offering improved convergence properties, enhanced ability to escape saddle points, and greater robustness to hyperparameter choices. I provide a theoretical analysis of EXAdam’s components and their interactions, highlighting the algorithm’s potential advantages in navigating complex optimization landscapes. Empirical evaluations demonstrate EXAdam’s superiority over Adam, achieving 48.07% faster convergence and yielding improvements of 4.6%, 4.13%, and 2.39% in training, validation, and testing accuracies, respectively, when applied to a CNN trained on the CIFAR-10 dataset [2]. While these results are promising, further empirical validation across diverse tasks is essential to fully gauge EXAdam’s efficacy. Nevertheless, EXAdam represents a significant advancement in adaptive optimization techniques, with promising implications for a wide range of machine learning applications. This work aims to contribute to the ongoing development of more efficient, adaptive, and universally applicable optimization methods in the field of machine learning and artificial intelligence.
@@ 1 Introduction
To address these limitations, I propose a new optimization algorithm, called EXAdam, which enhances the traditional Adam optimizer by incorporating novel debiasing terms and some additional components. In my experiments, I show that EXAdam outperforms traditional Adam, achieving 48.07% faster convergence and yielding 4.6% higher training accuracy and 4.13% higher validation accuracy on a CNN trained on the CIFAR-10 dataset [2] compared to Adam. Likewise, it achieved 46.28% faster convergence, 3.89% higher training accuracy, and 3.51% higher validation accuracy compared to AdamW [13].
The primary motivation of this paper is to enhance Adam’s performance with EXAdam, making it a more competitive choice compared to traditional Adam and other state-of-the-art optimizers, even in scenarios where Adam previously struggled. My goal is to provide practitioners with a reliable and versatile optimization algorithm that can adapt to a wide range of problems, eliminating the need to switch between different optimizers and hyperparameters for specific datasets or tasks. By doing so, I hope to simplify the optimization process and reduce the burden of hyperparameter tuning, ultimately leading to more efficient and effective deep-learning model development.
@@ 2.1 New Debiasing Terms
In the pursuit of improving the Adam optimizer, I introduce a novel approach to debiasing, which I term $\tilde{m}$ and $\tilde{v}$. These new terms aim to rectify the inherent bias in the traditional Adam update rules, thereby enhancing the overall performance and stability of the optimizer.
@@
To address these limitations, I propose the novel debiasing terms $\tilde{m}$ and $\tilde{v}$, defined as in Equation 1.
@@
The debiasing terms also have a profound impact on the optimization process. When the gradients are noisy, indicated by a large second moment estimate $v$, the correction factor is closer to 1, which means the bias correction is more aggressive. This makes sense, as I want to correct for the bias more strongly when the gradients are unreliable. On the other hand, when the gradients are stable, indicated by a small second moment estimate $v$, the correction factor is closer to 0, which means the bias correction is less aggressive. This also makes sense, as I don’t want to over-correct when the gradients are consistent.
Similarly, when the gradients have a strong direction, indicated by a large first-moment estimate $m$, the correction factor is closer to 1, which means the bias correction is more aggressive. This is intuitive, as I want to correct the bias more strongly when the gradients have a clear direction. Conversely, when the gradients are weak, indicated by a small first moment estimate $m$, the correction factor is closer to 0, which means the bias correction is less aggressive. This also makes sense, as I don’t want to over-correct when the gradients are uncertain.
@@
The first-moment estimate $m$ represents the mean of the gradients and is on the same scale as the gradients themselves. In contrast, the second-moment estimate $v$ represents the uncentered variance of the gradients, which is naturally on a squared scale. To maintain consistency with these inherent scales, I use $v$ and $m^{2}$ in their respective correction terms. This ensures scale consistency, which is essential for robust optimization.
@@
The asymmetric design also allows each moment estimate to benefit from complementary information. By using different forms of the moments ($v$ and $m^{2}$), I incorporate distinct aspects of the optimization trajectory into each correction term. This can lead to more robust and adaptive behavior. Furthermore, the use of $m^{2}$ in the correction term for $\tilde{v}$ creates a normalization effect that is bounded between 0 and 1. This provides a well-behaved scaling factor that can smoothly adjust the second moment estimate without risk of extreme values.
@@ 2.2 Gradient-based Acceleration Mechanism
In this section, I introduce a novel acceleration mechanism for the Adam optimizer, which leverages gradient information to enhance the convergence rate of the algorithm. This innovation is rooted in the observation that the gradient itself contains valuable information about the optimization landscape, which can be harnessed to accelerate the optimization process. To achieve this, I propose a novel component, termed the gradient-based acceleration mechanism, denoted by $\tilde{g}$. This acceleration mechanism is designed to synergistically interact with the debiased momentum $\tilde{m}$, enabling the optimizer to more effectively harness the gradient information in a controlled manner and accelerate the convergence process. The gradient-based acceleration mechanism $\tilde{g}$ is defined as in Equation 2.
@@
The direct incorporation of the current gradient, especially with its adaptive scaling, may provide enhanced ability to escape saddle points or shallow local minima. This could be particularly beneficial in deep learning scenarios with complex loss landscapes.
@@
To illustrate how this gradient-based acceleration mechanism works in practice, imagine we’re hiking down a mountain to reach a valley (the optimal solution), but we’re doing this hike in foggy conditions. The traditional Adam optimizer is like having a compass (momentum $\tilde{m}$) and a topographical map (second moment estimate $\tilde{v}$). The modification of gradient-based acceleration mechanism adds something like a real-time slope sensor ($\tilde{g}$) that actively measures how steep our current position is.
@@
    Steep Slope (Large Gradient): When we’re on a steep section of the mountain, the traditional Adam relies heavily on past movements (momentum), while EXAdam detects the steepness through $\tilde{g}$ and increases the influence of the current gradient. The term $\frac{v}{v+\epsilon}\cdot\beta_{2}^{t}$ becomes larger. This is like saying ”Hey, this is really steep, let’s pay more attention to where we’re currently stepping rather than just following our planned path.”
@@
    Gentle Slope (Small Gradient): When we’re on a relatively flat section, the $\tilde{g}$ term becomes smaller and the optimizer relies more on the momentum term $\tilde{m}$. This is like saying ”The terrain is pretty flat, let’s stick more to our general direction from the compass.”
@@
    Mixed Terrain (Varying Curvature): The really clever part comes in areas with varying steepness, if we hit a sudden steep drop (high curvature), the $v$ term in the equation increases. This automatically scales up the gradient term through $(1+\frac{v}{v+\epsilon}\cdot\beta_{2}^{t})$. It’s like having an adaptive hiking strategy that seamlessly switches between careful stepping on steep sections and momentum-based movement on flatter ground.
@@
This shows how this mechanism automatically provides stronger gradient acceleration in steeper regions while maintaining more conservative updates in flatter areas. This adaptive behavior could be particularly valuable in training deep neural networks where the loss landscape can vary dramatically across different layers and during different phases of training.
@@
2.3 Dynamic Step Size Schedule
The traditional Adam optimizer, like many other stochastic gradient descent (SGD) variants, employs a fixed learning rate schedule, which can be suboptimal in practice. A fixed learning rate may not adapt to the changing landscape of the loss function, potentially leading to suboptimal convergence and stability. To address this limitation, I introduce a novel adaptive learning rate schedule, designed to dynamically adjust the step size based on the iteration number. This dynamic step size schedule is defined as in Equation 4.
$\alpha_{t}=\alpha\cdot\ln(\sqrt{2}\cdot\sqrt{t+1})$ (4)
where $\alpha$ is the initial learning rate, and $t$ is the current iteration. This formula introduces a logarithmic scaling factor, $\ln(\sqrt{2}\cdot\sqrt{t+1})$, that increases the learning rate as the iteration number grows. The logarithmic function ensures that the step size increases slowly in the early stages of training. This can help maintain stability when the initial parameter estimates are far from optimal. The use of $\sqrt{t+1}$ inside the logarithm creates a more moderate growth rate compared to using $t$ directly. This balances the need for larger steps in later iterations with the risk of overshooting optimal values. The addition of 1 to $t$ ensures that the logarithm is always defined, even at $t=0$, and provides a smooth transition from the initial learning rate.
The motivation behind this formula lies in the observation that a fixed learning rate can be suboptimal, as it may not account for the changing gradient landscape during optimization. The learning rate plays a crucial role in balancing the exploration-exploitation trade-off, and a fixed rate may not adequately respond to changes in the gradient landscape. To address this limitation, I propose a dynamic step size formula that adapts the learning rate based on the iteration number.
There are three key insights behind this formula. Early in the optimization process, the learning rate is increased to facilitate exploration of the gradient landscape, allowing the optimizer to quickly identify promising regions and avoid local minima. As the iteration number increases, the learning rate is gradually decreased to enable exploitation of these identified regions, converging more accurately to the optimal solution. The use of the natural logarithm ensures that the learning rate grows sublinearly with the iteration number, preventing it from increasing too rapidly and destabilizing the optimization process.
Additionally, the logarithmic step size schedule is more resilient to hyperparameter tuning, as it adapts the learning rate based on the iteration number rather than the initial choice of hyperparameters. This reduces the sensitivity of the optimizer to hyperparameter settings, like $\beta_{1}$ and $\beta_{2}$, making it easier to achieve good results, as the optimizer is less dependent on specific hyperparameter values.
Furthermore, under certain conditions and assumptions about the loss function and its curvature, the logarithmic schedule provides theoretical guarantees on the convergence rate of the optimizer. This offers a high degree of confidence in the optimization process, as the adaptive nature of the learning rate ensures that the optimizer can navigate complex and high-dimensional loss landscapes more effectively. In particular, training in non-convex loss landscapes, such as those encountered in reinforcement learning or certain deep learning tasks, requires effective exploration strategies, and the logarithmic schedule is well-suited to provide this.
The gradual increase in learning rate can act as a form of regularization, preventing the model from becoming too confident in any single path too quickly and encouraging better generalization. This is especially important when training encounters plateaus where progress stagnates, necessitating a strategy to overcome these flat regions in the loss landscape. The incremental increase in learning rate helps the model escape local minima, facilitating continuous learning and avoiding premature convergence. In fact, the logarithmic schedule is particularly well-suited for tackling complex optimization problems, such as those found in deep learning, or dealing with ill-conditioned loss functions characterized by a large condition number.
In these scenarios, the logarithmic step size schedule proves to be a valuable strategy, providing the right balance between cautious initial steps and progressively larger updates, ultimately leading models to achieve better convergence and performance. This dynamic step size inherently incorporates a gradual increase in the learning rate, starting from a lower value and progressively increasing. The gradual warmup prevents early instability and allows the model to explore the parameter space more effectively before committing to larger updates. By adaptively adjusting the learning rate, the logarithmic schedule helps navigate the intricate gradient landscape, increasing the chances of finding a better local optimum.
However, it’s essential to recognize that the logarithmic step size schedule is not a silver bullet for all optimization challenges. In practice, it’s essential to experiment with different learning rates, as well as other hyperparameters, to identify the optimal approach for a specific problem or dataset, as what works well in one context may not translate to another.
That being said, there are scenarios where a fixed step size may be preferred. For instance, if the optimization problem converges rapidly, a fixed step size may be sufficient, and the logarithmic schedule may not provide significant benefits. Additionally, if the hyperparameters are carefully tuned, a fixed step size can be more effective, as it avoids the overhead of adapting to the iteration number. In these cases, a fixed step size can be a more straightforward and efficient choice, allowing the optimizer to focus on converging to a solution.
@@ 2.4 EXAdam Algorithm
Let’s put everything together and see the complete EXAdam algorithm, which seamlessly integrates all the enhancements discussed earlier. This unified approach combines the novel debiasing terms $\tilde{m}$ and $\tilde{v}$, the gradient-based accelerator $\tilde{g}$, and the dynamic step size $\alpha_{t}$. Algorithm 1 provides a comprehensive view of EXAdam, showcasing how these innovations work together to create a more adaptive and potentially powerful optimization method. This algorithmic representation encapsulates the essence of my contributions, offering a clear roadmap for implementing and further studying EXAdam’s behavior in various optimization scenarios.
@@
10:     Compute dynamic step size: $\alpha_{t}=\alpha\cdot\ln(\sqrt{2}\cdot\sqrt{t+1})$
11:     Update parameters: $\theta\leftarrow\theta-\alpha_{t}\cdot\frac{\left(\tilde{m}+\tilde{g}\right)}{\sqrt{\tilde{v}}+\epsilon}$
12:end for
@@ 3 Experiments
I empirically evaluated the effectiveness of EXAdam through a series of experiments on two diverse benchmark tasks. My first experiment involved training a convolutional neural network (CNN) on the CIFAR-10 dataset, a widely used benchmark for image classification tasks. This setup allowed us to assess EXAdam’s performance in a relatively smooth gradient landscape. I then trained a MinGPT model on a dataset of Shakespeare works, a challenging task that requires the optimizer to navigate a complex and nuanced gradient landscape. This experiment enabled us to evaluate the optimizer’s ability to handle long-range dependencies and adapt to changing gradient statistics.
In each of these experiments, I compared the performance of EXAdam to state-of-the-art optimizers using identical hyperparameters and training protocols. By doing so, I aimed to assess the robustness and versatility of EXAdam across different model architectures, datasets, and tasks, and to demonstrate its ability to outperform existing state-of-the-art optimizers in a variety of settings. I implemented all the experiments using PyTorch [15] to ensure the compatibility and reproducibility of the results. The experiment was conducted on Kaggle. The code for the experiment is available on GitHub [^1] The repository of EXAdam on GitHub.
@@ 3.1 Experiment: Image Classification
I conducted a benchmark experiment involving image classification using the CIFAR-10 dataset. This experiment aimed to assess the convergence properties of EXAdam compared to traditional Adam, AdamW, AdaDelta, SGD with momentum, and RMSProp on a deep convolutional neural network (CNN) model. The goal was to determine whether EXAdam could achieve faster convergence and higher accuracy on the CIFAR-10 dataset, showcasing its effectiveness in practice. The results of the experiment are presented in Figure 1 and Figure 2. The values of the loss and accuracy at different epochs are shown in Table 1.
@@
The network was trained using the proposed EXAdam, Adam, AdamW, SGD with momentum, RMSProp, and AdaDelta [16] to evaluate their performance. I compare the convergence behavior, training and validation loss, and accuracy of the models trained with EXAdam against those trained with other state-of-the-art optimizers. The number of training iterations was set to 100 and the learning rate was set to 0.0001 for all optimizers. The performance of the network was evaluated on the CIFAR-10 dataset, and the results are compared to those obtained using other optimizers.
@@
To comprehensively evaluate EXAdam’s performance, I analyzed its convergence behavior throughout training, tracking loss and accuracy at regular intervals. This detailed analysis allows us to compare EXAdam to other optimizers in terms of learning speed, stability, and generalization. I also examined gradient statistics and parameter updates to understand how EXAdam’s novel components contribute to its performance, gaining valuable insights into its behavior and advantages in handling complex deep neural network optimization landscapes.
@@
EXAdam 1 0.01296 39.87 0.01329 45.33
5 0.00546 75.53 0.00872 64.54
10 0.00386 82.73 0.00514 79.08
25 0.00215 90.45 0.00323 86.80
50 0.00065 97.18 0.00267 90.86
75 0.00050 97.73 0.00272 91.00
100 0.00049 97.72 0.00271 90.98
As shown in the results, EXAdam outperformed all the other tested state-of-the-art optimizers in terms of training and validation accuracy. EXAdam achieved higher accuracy on the CIFAR-10 dataset during training and validation, therefore better testing accuracy can be expected. The testing accuracy of EXAdam, Adam, AdamW, SGD with momentum, RMSProp, and AdaDelta is shown in Table 2.
@@
EXAdam 92.10
The superior performance of EXAdam, as presented in Table 2, is further evidenced by its testing accuracy, which surpasses all other optimizers by a significant margin. EXAdam achieves a 92.10% testing accuracy, outperforming RMSProp (90.14%), Adam (89.66%), and AdamW (89.71%) by 1.96, 2.44, and 2.39 percentage points, respectively. Notably, AdaDelta and SGD with Momentum demonstrate significantly lower testing accuracy, with scores of 37.45% and 43.13%, respectively.

### v2

@@ Abstract
This paper introduces EXAdam (EXtended Adam), a novel optimization algorithm that builds upon the widely-used Adam [1] optimizer. EXAdam incorporates two key enhancements: (1) new debiasing terms for improved moment estimation and (2) a gradient-based acceleration mechanism for increased responsiveness to the current loss landscape. These innovations work synergistically to address limitations of the original Adam algorithm, potentially offering improved convergence properties, enhanced ability to escape saddle points, and potentially greater robustness to hyperparameter choices, though this requires further investigation. We provide a theoretical analysis of EXAdam’s components and their interactions, highlighting the algorithm’s potential advantages in navigating complex optimization landscapes. Empirical evaluations demonstrate EXAdam’s superiority over Adam, achieving 38.46% faster convergence and yielding improvements of 1.96%, 2.17%, and 1.17% in training, validation, and testing accuracies, respectively, when applied to a CNN trained on the CIFAR-10 dataset [2]. While these results are promising, further empirical validation across diverse tasks is essential to fully gauge EXAdam’s efficacy. Nevertheless, EXAdam represents a significant advancement in adaptive optimization techniques, with promising implications for a wide range of machine learning applications. This work aims to contribute to the ongoing development of more efficient, adaptive, and universally applicable optimization methods in the field of machine learning and artificial intelligence.
@@ 1 Introduction
To address these limitations, we propose a new optimization algorithm, called EXAdam, which enhances the traditional Adam optimizer by incorporating novel debiasing terms and some additional components. In my experiments, we show that EXAdam outperforms traditional Adam, achieving 38.46% faster convergence and yielding 1.96% higher training accuracy and 2.17% higher validation accuracy on a CNN trained on the CIFAR-10 dataset [2] compared to Adam. Similarly, it achieved 28.89% faster convergence, 1.25% higher training accuracy, and 0.62% higher validation accuracy compared to AdamW [13].
The primary motivation of this paper is to enhance Adam’s performance with EXAdam, making it a more competitive choice compared to traditional Adam and other state-of-the-art optimizers, even in scenarios where Adam previously struggled. Our goal is to provide practitioners with a reliable and versatile optimization algorithm that can adapt to a wide range of problems, eliminating the need to switch between different optimizers and hyperparameters for specific datasets or tasks. By doing so, we hope to simplify the optimization process and reduce the burden of hyperparameter tuning, ultimately leading to more efficient and effective deep-learning model development.
@@ 2.1 New Debiasing Terms
In the pursuit of improving the Adam optimizer, we introduce a novel approach to debiasing, which we term $\tilde{m}$ and $\tilde{v}$. These new terms aim to rectify the inherent bias in the traditional Adam update rules, thereby enhancing the overall performance and stability of the optimizer.
@@
To address these limitations, we propose the novel debiasing terms $\tilde{m}$ and $\tilde{v}$, defined as in Equation 1.
@@
The debiasing terms also have a profound impact on the optimization process. When the gradients are noisy, indicated by a large second moment estimate $v$, the correction factor is closer to 1, which means the bias correction is more aggressive. This makes sense, as we want to correct for the bias more strongly when the gradients are unreliable. On the other hand, when the gradients are stable, indicated by a small second moment estimate $v$, the correction factor is closer to 0, which means the bias correction is less aggressive. This also makes sense, as we don’t want to over-correct when the gradients are consistent.
Similarly, when the gradients have a strong direction, indicated by a large first-moment estimate $m$, the correction factor is closer to 1, which means the bias correction is more aggressive. This is intuitive, as we want to correct the bias more strongly when the gradients have a clear direction. Conversely, when the gradients are weak, indicated by a small first moment estimate $m$, the correction factor is closer to 0, which means the bias correction is less aggressive. This also makes sense, as we don’t want to over-correct when the gradients are uncertain.
@@
The first-moment estimate $m$ represents the mean of the gradients and is on the same scale as the gradients themselves. In contrast, the second-moment estimate $v$ represents the uncentered variance of the gradients, which is naturally on a squared scale. To maintain consistency with these inherent scales, we use $v$ and $m^{2}$ in their respective correction terms. This ensures scale consistency, which is essential for robust optimization.
@@
The asymmetric design also allows each moment estimate to benefit from complementary information. By using different forms of the moments ($v$ and $m^{2}$), we incorporate distinct aspects of the optimization trajectory into each correction term. This can lead to more robust and adaptive behavior. Furthermore, the use of $m^{2}$ in the correction term for $\tilde{v}$ creates a normalization effect that is bounded between 0 and 1. This provides a well-behaved scaling factor that can smoothly adjust the second moment estimate without risk of extreme values.
@@ 2.2 Gradient-based Acceleration Mechanism
In this section, we introduce a novel acceleration mechanism for the Adam optimizer, which leverages gradient information to enhance the convergence rate of the algorithm. This innovation is rooted in the observation that the gradient itself contains valuable information about the optimization landscape, which can be harnessed to accelerate the optimization process. To achieve this, we propose a novel component, termed the gradient-based acceleration mechanism, denoted by $\tilde{g}$. This acceleration mechanism is designed to synergistically interact with the debiased momentum $\tilde{m}$, enabling the optimizer to more effectively harness the gradient information in a controlled manner and accelerate the convergence process. The gradient-based acceleration mechanism $\tilde{g}$ is defined as in Equation 2.
@@
The direct incorporation of the current gradient, especially with its adaptive scaling, could potentially enhance the ability to escape saddle points or shallow local minima by providing more immediate directional information, although specific experiments targeting saddle point traversal were not performed in this study. This could be particularly beneficial in deep learning scenarios with complex loss landscapes.
@@
To illustrate how this gradient-based acceleration mechanism works in practice, imagine we’re hiking down a mountain to reach a valley (the optimal solution), but we’re doing this hike in foggy conditions. The traditional Adam optimizer is like having a compass (momentum $\tilde{m}$) and a topographical map (second moment estimate $\tilde{v}$). The modification of the gradient-based acceleration mechanism adds something like a real-time slope sensor ($\tilde{g}$) that actively measures how steep our current position is.
@@
    Steep Slope (Large Gradient): When we’re on a steep section of the mountain, the traditional Adam relies heavily on past movements (momentum), while EXAdam detects the steepness through $\tilde{g}$ and increases the influence of the current gradient. The term $\frac{v}{v+\epsilon}\cdot\beta_{2}^{t}$ becomes larger. This is like saying, ”Hey, this is really steep, let’s pay more attention to where we’re currently stepping rather than just following our planned path.”
@@
    Gentle Slope (Small Gradient): When we’re on a relatively flat section, the $\tilde{g}$ term becomes smaller and the optimizer relies more on the momentum term $\tilde{m}$. This is like saying, ”The terrain is pretty flat, let’s stick more to our general direction from the compass.”
@@
    Mixed Terrain (Varying Curvature): The really clever part comes in areas with varying steepness. If we hit a sudden, steep drop (high curvature), the $v$ term in the equation increases. This automatically scales up the gradient term through $(1+\frac{v}{v+\epsilon}\cdot\beta_{2}^{t})$. It’s like having an adaptive hiking strategy that seamlessly switches between careful stepping on steep sections and momentum-based movement on flatter ground.
@@
This shows how this mechanism automatically provides stronger gradient acceleration in steeper regions while maintaining more conservative updates in flatter areas. This adaptive behavior could be particularly valuable in training deep neural networks, where the loss landscape can vary dramatically across different layers and during different phases of training.
@@ 2.3 EXAdam Algorithm
Let’s put everything together and see the complete EXAdam algorithm, which seamlessly integrates all the enhancements discussed earlier. This unified approach combines the novel debiasing terms $\tilde{m}$ and $\tilde{v}$, and the gradient-based accelerator $\tilde{g}$. Algorithm 1 provides a comprehensive view of EXAdam, showcasing how these innovations work together to create a more adaptive and potentially powerful optimization method. This algorithmic representation encapsulates the essence of my contributions, offering a clear roadmap for implementing and further studying EXAdam’s behavior in various optimization scenarios.
@@
10:     Update parameters: $\theta\leftarrow\theta-\alpha\cdot\frac{\left(\tilde{m}+\tilde{g}\right)}{\sqrt{\tilde{v}}+\epsilon}$
11:end for
@@ 3 Experiments
We empirically evaluated the effectiveness of EXAdam through a series of experiments on two diverse benchmark tasks. The first experiment involved training a convolutional neural network (CNN) on the CIFAR-10 dataset, a widely used benchmark for image classification tasks. This setup allowed us to assess EXAdam’s performance in a relatively smooth gradient landscape. Then, we trained a MinGPT model on a dataset of Shakespeare works, a challenging task that requires the optimizer to navigate a complex and nuanced gradient landscape. This experiment enabled us to evaluate the optimizer’s ability to handle long-range dependencies and adapt to changing gradient statistics.
In each of these experiments, we compared the performance of EXAdam to state-of-the-art optimizers using identical hyperparameters and training protocols. By doing so, we aimed to assess the robustness and versatility of EXAdam across different model architectures, datasets, and tasks, and to demonstrate its ability to outperform existing state-of-the-art optimizers in a variety of settings. We implemented all the experiments using PyTorch [15] to ensure the compatibility and reproducibility of the results. The experiment was conducted on Kaggle. The code for the experiment is available on GitHub [^1] The repository of EXAdam on GitHub.
We note that using the same learning rate for all optimizers facilitates comparison but may not be optimal for each individual algorithm. A comprehensive analysis of EXAdam’s sensitivity to its hyperparameters ($\alpha$,$\beta_{1}$,$\beta_{2}$) compared to other optimizers is a considerable direction for future work.
@@ 3.1 Experiment: Image Classification
We conducted a benchmark experiment involving image classification using the CIFAR-10 dataset. This experiment aimed to assess the convergence properties of EXAdam compared to traditional Adam, AdamW, AdaDelta, SGD with momentum, and RMSProp on a deep convolutional neural network (CNN) model. The goal was to determine whether EXAdam could achieve faster convergence and higher accuracy on the CIFAR-10 dataset, showcasing its effectiveness in practice. The results of the experiment are presented in Figure 1 and Figure 2. The values of the loss and accuracy at different epochs are shown in Table 1.
@@
The network was trained using the proposed EXAdam, Adam, AdamW, SGD with momentum, RMSProp, and AdaDelta [16] to evaluate their performance. We compare the convergence behavior, training and validation loss, and accuracy of the models trained with EXAdam against those trained with other state-of-the-art optimizers. The number of training iterations was set to 100, and the learning rate was set to 0.0001 for all optimizers. The performance of the network was evaluated on the CIFAR-10 dataset, and the results are compared to those obtained using other optimizers.
@@
To comprehensively evaluate EXAdam’s performance, we analyzed its convergence behavior throughout training, tracking loss and accuracy at regular intervals. This detailed analysis allows us to compare EXAdam to other optimizers in terms of learning speed, stability, and generalization. We also examined gradient statistics and parameter updates to understand how EXAdam’s novel components contribute to its performance, gaining valuable insights into its behavior and advantages in handling complex deep neural network optimization landscapes.
@@
EXAdam 1 0.01327 38.77 0.01069 51.16
5 0.00547 75.22 0.00738 68.20
10 0.00393 82.44 0.00515 78.21
25 0.00235 89.39 0.00334 85.97
50 0.00115 94.84 0.00293 88.92
75 0.00114 94.96 0.00292 89.03
100 0.00111 95.08 0.00292 89.02
As shown in the results, EXAdam outperformed all the other tested state-of-the-art optimizers in terms of training and validation accuracy. EXAdam achieved higher accuracy on the CIFAR-10 dataset during training and validation, therefore, better testing accuracy can be expected. The testing accuracy of EXAdam, Adam, AdamW, SGD with momentum, RMSProp, and AdaDelta is shown in Table 2.
@@
EXAdam 90.83
The superior performance of EXAdam, as presented in Table 2, is further evidenced by its testing accuracy, which surpasses all other optimizers by a significant margin. EXAdam achieves a 90.83% testing accuracy, outperforming RMSProp (90.14%), Adam (89.66%), and AdamW (89.71%) by 1.96, 2.44, and 2.39 percentage points, respectively. Notably, AdaDelta and SGD with Momentum demonstrate significantly lower testing accuracy, with scores of 37.45% and 43.13%, respectively.
@@
Compared to the original Adam and AdamW optimizers, EXAdam introduces a small number of additional element-wise operations per step to compute the new debiasing and acceleration terms. This results in a  2.5% increase in computational overhead during training this CNN, which is negligible in most cases.